OpenStudy (anonymous):

find the sin A and cos A. write your ans as a fraction and as decimals rounded to four decimal places

OpenStudy (anonymous):



OpenStudy (campbell_st):

so your reference angle is A  \[Sin A =\frac{opp}{hyp}....and..cosA =\frac{adj}{hyp}\] \[\sin(A) = \frac{\sqrt{11}}{6}\] and \[\cos(A) = \frac{5}{6}\] hope this helps